recently discovered and found to have an automatic weight of 342.38


1
Expert's answer
2021-10-24T01:28:16-0400

Atomic weight of compound =342.38am

Mass of lighter isotope= 340.91amu

"\%" abundance of lighter isotope

=68.322"\%"

Mass of lighter isotope in the compound

"= \frac{68.322}{100}x 340.91"

"= 232.9165amu"

Mass of heavier isotope in the compound

=342.38 - 232.9165

=109.4635amu

"\%" abundance of heavier isotope

="100\%-68.322\%"

=31.678"\%"


"\therefore\frac {31.678}{100}×" mass of the heavier isotope

= 109.4635

"\implies" mass of heavier isotope

"\frac {109.4635}{31.678}×100"


= 345.55 amu